# Hierarchical Clustering for Conditional Diffusion in Image Generation

### v1

@@ 4 Experiments
We present a series of experiments designed to evaluate the performance of TreeDiffusion across various datasets. In Section 4.1, we compare the clustering and generative performance of TreeDiffusion to TreeVAE (Manduchi et al., 2023) using several benchmark datasets, including MNIST (Lecun et al., 1998), FashionMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky, 2009), and CUBICC (Palumbo et al., 2023). The CUBICC dataset, a variant of the CUB Image-Captions dataset (Wah et al., 2011; Shi et al., 2019), contains images of birds grouped into eight specific species, allowing for a detailed analysis of clustering performance. Additionally, we conduct generative evaluations using the CelebA dataset (Liu et al., 2015) to assess the model’s ability to generate high-quality images. In Section 4.2, we assess how cluster-specific the generated images are and analyze the variability among samples generated from the same cluster, examining whether there are any indications of mode collapse. Finally, in Section 4.3, we perform an ablation study on the conditioning signals to compare the generative capabilities of different model configurations and identify the signals that most effectively enhance performance. Through these experiments, we aim to demonstrate the effectiveness of the TreeDiffusion model in both clustering and generation tasks.
4.1 Generative and Clustering Performance
@@
The following analysis compares two models: the TreeVAE and the proposed TreeDiffusion. The TreeDiffusion models are conditioned on the path $\bm{z}_{\mathcal{P}_{l}}$ retrieved from the TreeVAE. Reconstruction performance is assessed using the Fréchet Inception Distance (Heusel et al., 2017, FID), calculated for the reconstructed images from the images in the test set. Additionally, we compute the FID score using $10,000$ newly generated images to evaluate generative performance. Clustering performance is measured using accuracy (ACC) and normalized mutual information (NMI). Table 1 presents the results of this analysis.
Table 1: Test set generative and clustering performances of different TreeVAE models. Means and standard deviations are computed across 10 runs with different seeds.
Dataset Method FID (rec) $\downarrow$ FID (gen) $\downarrow$ ACC $\uparrow$ NMI $\uparrow$
[…]
TreeDiffusion 1.5± $\pm\,0.0$ 1.8± $\pm\,0.1$
[…]
TreeDiffusion 5.5± $\pm\,0.6$ 5.4± $\pm\,0.4$
[…]
TreeDiffusion 12.5± $\pm\,0.4$ 17.8± $\pm\,0.4$
[…]
TreeDiffusion 13.4± $\pm\,0.9$ 29.0± $\pm\,5.4$
[…]
@@
Notably, the clustering performance remains identical for both models, as TreeDiffusion leverages the hierarchical clustering information from the pre-trained TreeVAE. The results indicate that TreeDiffusion significantly enhances the generative capabilities of the model, reducing FID scores by roughly an order of magnitude across all datasets. This improvement is particularly evident in the quality of the generated images, as illustrated in Figure 2 for the CUBICC dataset. Here, we visually compare the samples generated by TreeDiffusion with those produced by the underlying TreeVAE, which supplied the cluster path as the conditioning signal for TreeDiffusion. The TreeVAE model continues to produce visibly blurry images, whereas TreeDiffusion generates noticeably sharper samples that adhere better to the data distribution.
Figure 2: (Top) Ten different samples generated by the TreeVAE model, each generated by sampling one path in the tree. (Bottom) Corresponding samples from the TreeDiffusion model, conditioned on the same selected path and embeddings from TreeVAE.
Figure 3: Image generations from every leaf of the (top) TreeVAE and (bottom) TreeDiffusion model, both trained on the CUBICC dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample. The corresponding leaf probabilities are shown above each image and are identical across both models by design.
@@ 4.2 Cluster-specific Representations
Higher quality cluster-specific generations
In Figure 2, we present randomly generated images for the CUBICC dataset for both TreeVAE and TreeDiffusion, where each column corresponds to one generation process. For each generation, we first sample the root embedding; then, we sample the path in the tree and the refined representations along the selected path iteratively until a leaf is reached. The hierarchical representation is then used to condition the inference in TreeDiffusion. As can be seen, the TreeDiffusion generations show substantially higher generative quality. Additionally, we examine further the first generated sample from Figure 2. For this sample, we present the generations from all leaves in Figure 3 by propagating the corresponding root representation across all paths in the tree. Note that leaf ”L2” has the highest path probability across all leaves. When comparing the generated images across the leaves for both models, it is evident that TreeDiffusion not only produces sharper images for all clusters but also generates a greater diversity of images. We ensured the same level of stochasticity for both models, eliminating potential confounding factors. Therefore, the observed diversity stems from the models themselves. As a result, the TreeDiffusion model can generate cluster-specific images that preserve the overall color and structure seen in TreeVAE images while significantly enhancing the distinctiveness and clarity of the images for each cluster. Further examples of the leaf-specific image generations can for TreeVAE and TreeDiffusion can be found in C.2.
@@
Figure 5: TreeDiffusion model trained on CIFAR-10. For each cluster, random newly generated images are displayed. Below each set of images, a normalized histogram (ranging from 0 to 1) shows the distribution of predicted classes from an independent, pre-trained classifier on CIFAR-10 for all newly generated images in each leaf with a significant probability of reaching that leaf.
@@
Figure 6: Image generations from each leaf of (top) a TreeVAE, (middle) a DiffuseVAE which only conditions on the reconstruction from the TreeVAE, and (bottom) a TreeDiffusion model conditioned on the path embeddings, all trained on CUBICC. Each row displays the generated images from all leaves of the specified model, starting with the same sample from the root. The corresponding leaf probabilities are shown at the top of the image and are, by design, the same for all models.
Hierarchical information is retained across generations
@@
To assess whether the newly generated images retain their hierarchical information, we train a classifier on the original training datasets and then utilize it to classify the newly generated images from our TreeDiffusion. Specifically, we classify the newly generated images for each cluster separately. Ideally, “pure“ leaves should be characterized by leaf-generations that are classified into one or very few classes from the original dataset. For this classification task, we utilize a ResNet-50 model (He et al., 2016) trained on each dataset. In Figure 4, we present randomly generated images from a TreeDiffusion model trained on FashionMNIST, together with normalized histograms depicting the distribution of the predicted classes for each leaf. For instance, clusters representing trousers and bags appear to accurately and distinctly capture their respective classes, as all their generated images are classified into one group only. Conversely, certain clusters are characterized by a mixture of classes, indicating that they are grouped together. Further results can be observed for the CIFAR-10 or MNIST dataset, shown in Figure 5 and Figure 8, respectively. Overall, we observe that the leaf-specific generations retain the hierarchical clustering structure found by TreeVAE, thereby enhancing interpretability in diffusion models.
On the benefits of hierarchical conditioning
@@
We hereby assess whether the conditioning on hierarchical representations improves cluster-specific generative quality. To this end, we compare the generations of TreeDiffusion, which is conditioned on the hierarchical representation, to a baseline, here defined cluster-unconditional, that is conditioned only on the leaf-reconstructions. For this experiment, we use the previously introduced independent classifier to create histograms for each leaf to evaluate how cluster-specific the newly generated images are. As previously mentioned, ideally, the majority of generated images from one leaf should be classified into one or very few classes from the original dataset. To quantify this, we compute the average entropy for all leaf-specific histograms. Lower entropy indicates less variation in the histograms and, thus, more leaf-specific generations.
@@
Table 2: Cluster-specificity of TreeDiffusion generations for cluster-unconditional and cluster-conditional reverse models, measured by mean entropy. Lower entropy indicates more cluster-specific generations. The best result for each dataset is marked in bold.
Dataset Method Mean Entropy
MNIST unconditional $1.24$
conditional 0.33
Fashion unconditional $0.66$
conditional 0.65
CIFAR10 unconditional $1.12$
conditional 0.93
CUBICC unconditional 0.07
conditional $0.20$
@@
Table 2 presents the results across all datasets. The conditional model consistently shows lower mean entropy, indicating that, for most datasets, conditioning on the hierarchy indeed helps guide the model to generate more distinct and representative images for each leaf. However, for the CUBICC dataset, we observe that the mean entropy is lower for the cluster-unconditional model. This is because the classifier tends to predict all images into a single class, a result of model degeneration, where it primarily generates images for only a few classes. Figure 6 visually presents the leaf generations for one sample of these models alongside the underlying TreeVAE generations. It can be observed that both the cluster-unconditional and conditional models exhibit a significant improvement in image quality. However, the images in the cluster-conditional model are more diverse, demonstrating greater adaptability for each cluster. Notably, across all models, the leaf-specific images share common properties, such as background color and overall shape, sampled at the root while varying in cluster-specific features from leaf to leaf within each model.
@@ 4.3 Ablation study on conditioning information
Finally, we perform an ablation study to assess the effects of various conditioning signals on the generative performance of the proposed approach. The results, outlined in Table 3, show the FID score calculated from $10,000$ samples generated using $100$ DDIM steps, averaged over $10$ random seeds. The findings indicate that utilizing information from the latent leaf — whether through leaf assignment, leaf embedding, or both — yields better generative performance compared to using only the leaf reconstruction. Additionally, conditioning on the full path $\bm{z}_{\mathcal{P}_{l}}$, which incorporates all embeddings and intermediate node assignments from the root to the leaf, further enhances the performance, underscoring the effectiveness of hierarchical clustering information. Notably, conditioning on the latent path also exceeds the generative performance of a fully unconditional DDIM model, which achieves an average FID of $18.1$. Consequently, utilizing the path $\bm{z}_{\mathcal{P}_{l}}$ from the hierarchical structure not only results in more structured generations, as demonstrated in Figure 3, but can also enhance the generative performance of generative clustering models.
Table 3: Effect of conditioning signals on generative performance for CIFAR-10. FID scores for $00010^{\prime}$ 000 samples (lower is better) computed across $10$ random model initializations.
Leaf Reconstruction Leaf Assignment Leaf Embedding Path FID $\downarrow$
$\hat{{\bm{x}}}_{0}^{(l)}$ $l$ $\bm{z}_{l}$ $\bm{z}_{\mathcal{P}_{l}}$
[…]
@@ 5 Conclusion
In this work, we present TreeDiffusion, a novel approach to integrate hierarchical clustering into diffusion models. By enhancing TreeVAE with a Denoising Diffusion Implicit Model conditioned on cluster-specific representations, we propose a model capable of generating distinct, high-quality images that faithfully represent their respective data clusters. This approach not only improves the visual fidelity of generated images but also ensures that these representations are true to the underlying data distribution. TreeDiffusion offers a robust framework that bridges the gap between clustering precision and generative performance, thereby expanding the potential applications of generative models in areas requiring detailed and accurate visual data interpretation.
Acknowledgments and Disclosure of Funding
@@
Laura Manduchi is supported by the SDSC PhD Fellowship #1-001568-037. Moritz Vandenhirtz is supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00047.
Appendix A Detailed Formulation of TreeVAE

### v2

@@ 4 Experiments
We present a series of experiments to evaluate the performance of TreeDiffusion across various datasets. The experiments are carried out on MNIST [16], FashionMNIST [42], CIFAR-10 [15], CelebA [18], and CUBICC [22]. The CUBICC dataset is a variant of the CUB Image-Captions dataset [36, 29], consisting of bird images grouped into eight distinct species. In Section 4.1, we compare the generative performance of TreeDiffusion against baseline methods. In Section 4.2, we evaluate how specific the generated images are to their source clusters and analyze how distinct images from different clusters are from one another. Finally, in Section 4.3, we perform an ablation study on conditioning signals, examining various model configurations to identify which signals most effectively improve generative performance. Specifically, we compare conditioning strategies based on hierarchical clustering, flat clustering, and no clustering information.
4.1 Generative Performance
@@
The following analysis compares the proposed TreeDiffusion model with TreeVAE [19], and a naive hybrid approach — referred to as "TreeVAE + Diffusion". In this hybrid approach, the diffusion model refines the TreeVAE output and refines it but without being conditioned on any latent information, as it is done in DiffuseVAE [23]. In contrast, TreeDiffusion conditions the diffusion model on the path information of the selected cluster retrieved from TreeVAE. For both these models, the TreeVAE instances serve as the first-stage model. The evaluation considers both reconstruction and generative performance, measured using the Fréchet Inception Distance (FID) [10]. Reconstruction performance is assessed by computing the FID score between reconstructed images and their corresponding test set images. Generative performance is evaluated by calculating the FID score for 10,000 newly generated images. The results of this analysis are summarized in Table 1.
The naive approach (TreeVAE + Diffusion) and TreeDiffusion both achieve substantial improvements over the baseline TreeVAE, their first-stage model, reducing FID scores by approximately an order of magnitude across all datasets. The naive approach performs better on simpler grayscale datasets, primarily excelling at image reconstruction rather than generation. This highlights a tendency toward overfitting. Conversely, TreeDiffusion consistently outperforms on the more complex, real-world color datasets at generating new images. Most likely, the difference in performance stems from how the denoising process is initialized. The naive model begins denoising from TreeVAE reconstructions, thereby making it highly dependent on the reconstruction quality provided by TreeVAE. Given that TreeVAE struggles more with generating new images than with reconstruction, this limitation is propagated into the naive approach. TreeDiffusion circumvents this issue by initializing the denoising directly from noise, using only latent representations from TreeVAE. As a result, TreeDiffusion achieves a better balance between reconstruction and generation quality, leading to better FID scores on newly generated images. Figure˜2 compares image reconstructions on CIFAR-10, demonstrating that both diffusion-based models significantly improve upon the image quality produced by TreeVAE.
Table 1: Test set generative performances of the compared models. FID scores for $10,000$ samples (lower is better) computed across $10$ random model initializations.
Dataset Method FID (rec) $\downarrow$ FID (gen) $\downarrow$
MNIST TreeVAE $24.0\,\text{{\footnotesize$\pm\,0.9$}}$ $21.8\,\text{{\footnotesize$\pm\,0.7$}}$
TreeVAE + Diffusion 1.4± $\pm\,0.0$ 1.8± $\pm\,0.1$
TreeDiffusion $1.5\,\text{{\footnotesize$\pm\,0.0$}}$ 1.8± $\pm\,0.1$
Fashion TreeVAE $40.7\,\text{{\footnotesize$\pm\,2.1$}}$ $41.9\,\text{{\footnotesize$\pm\,2.1$}}$
TreeVAE + Diffusion 4.8± $\pm\,0.2$ 4.8± $\pm\,0.2$
TreeDiffusion 5.5± $5\,\text{{\footnotesize$\pm\,0.6$}}$ $5.4\,\text{{\footnotesize$\pm\,0.4$}}$
CIFAR-10 TreeVAE $175.8\,\text{{\footnotesize$\pm\,1.4$}}$ $188.0\,\text{{\footnotesize$\pm\,2.0$}}$
TreeVAE + Diffusion 12.3± $\pm\,0.1$ $19.7\,\text{{\footnotesize$\pm\,0.2$}}$
TreeDiffusion $12.5\,\text{{\footnotesize$\pm\,0.4$}}$ 17.8± $\pm\,0.4$
CUBICC TreeVAE $232.5\,\text{{\footnotesize$\pm\,7.1$}}$ $255.3\,\text{{\footnotesize$\pm\,8.8$}}$
TreeVAE + Diffusion 12.7± $\pm\,6.5$ $96.0\,\text{{\footnotesize$\pm\,2.1$}}$
TreeDiffusion $13.4\,\text{{\footnotesize$\pm\,0.9$}}$ 29.0± $\pm\,5.4$
CelebA TreeVAE $5.2\,\text{{\footnotesize$\pm\,15.0$}} 77$$7.9\,\text{{\footnotesize$\pm\,5.6$}}$
TreeVAE + Diffusion $15.4\,\text{{\footnotesize$\pm\,3.2$}}$ $30.1\,\text{{\footnotesize$\pm\,7.5$}}$
@@
Figure 2: Ten different CIFAR-10 reconstructions generated by the TreeVAE model, each obtained by sampling a single path in the tree. Corresponding reconstructions from TreeVAE + Diffusion, which begins denoising with the TreeVAE reconstructions, are shown alongside those from TreeDiffusion, which conditions on the same selected path and embeddings but starts denoising from noise.
Figure 3: Ten different samples generated by the TreeVAE model, each generated by sampling one path in the tree, and corresponding samples from the TreeDiffusion model, conditioned on the same selected path and embeddings from TreeVAE.
Figure 4: Image generations from every leaf of the TreeVAE and TreeDiffusion model, both trained on the CUBICC dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample.
@@ 4.2 Cluster-specific Representations
4.2.1 Higher quality cluster-specific generations.
In Figure˜3, we present randomly generated images for the CUBICC dataset for both TreeVAE and TreeDiffusion, where each column corresponds to an independently generated sample. For each generation, we first sample the root embedding; then, we sample the path in the tree and the refined representations along the selected path iteratively until a leaf is reached. The hierarchical representation is then used to condition the inference in TreeDiffusion. As can be seen, the TreeDiffusion generations show substantially higher generative quality. In the following, we examine the first generated sample from Figure˜3 in more detail. For this one sample, we present the generations of all leaves in Figure˜4 by propagating the corresponding root representation across all paths in the tree. Note that the selected sample shown in Figure˜3 ended up stemming from leaf 3 in Figure˜4. When comparing the generated images across the leaves for both models, it is evident that TreeDiffusion not only produces sharper images for all clusters but also generates a greater diversity of images. Note that both models utilize the same latent information for image generation. While TreeVAE and TreeDiffusion maintain similar overall color distribution and structural characteristics, TreeDiffusion significantly enhances cluster specificity, resulting in images with greater clarity and distinctiveness for each cluster.
@@
Figure 6: Image generations from each leaf of (top) TreeVAE, (middle) TreeVAE + Diffusion which starts denoising with the TreeVAE images, and (bottom) TreeDiffusion model conditioned on the hierarchical path embeddings, all trained on CUBICC. Each row displays the generated images from all leaves of the specified model, starting with the same sample from the root. The corresponding leaf probabilities are shown at the top of the image and are, by design, the same for all models.
4.2.2 Cluster information is retained across generations.
@@
To quantitively assess whether the newly generated images retain their cluster information, we train a classifier on the original labeled training data and then use it to classify generated images of TreeDiffusion. Specifically, we classify the generations for each cluster separately. The idea is that “pure“ leaves should create samples that are classified into one or very few classes. For this classification task, we use a ResNet-50 model [9] trained on each dataset. In Figure˜5, we present randomly generated images from a TreeDiffusion model trained on FashionMNIST, together with normalized histograms depicting the distribution of the predicted classes for each leaf. For instance, clusters representing trousers and bags appear to accurately and distinctly capture their respective classes, as all their generated images are classified into one group only. Conversely, certain clusters are characterized by a mixture of classes, indicating that they are grouped together. Overall, we observe that the leaf-specific generations retain the hierarchical clustering structure found by TreeVAE, thereby enhancing the interpretability in diffusion models.
4.2.3 On the benefits of hierarchical conditioning.
@@
We hereby assess whether the conditioning on hierarchical representations improves cluster-specific generative quality. To this end, we compare the generations of TreeDiffusion, which is conditioned on the hierarchical representation, to the baseline TreeVAE + Diffusion from earlier, which is not conditioned on the latent cluster information. For this experiment, we use the previously introduced independent classifier to create the normalized histograms for each leaf to evaluate how cluster-specific the newly generated images are. As mentioned above, ideally, the majority of generated images from one leaf should be classified into one or very few classes from the original dataset. To quantify this, we compute the average entropy for all cluster-specific histograms. Lower entropy indicates less variation in the histograms and, thus, more cluster-specific generations. Table 2 presents the results for all labeled datasets.
@@
Table 2: Cluster-specificity of TreeDiffusion generations comparing cluster-unconditional and cluster-conditional reverse models, measured by mean entropy. Lower entropy indicates more cluster-specific generations. The best result for each dataset is marked in bold.
Dataset Method Cluster Conditioning Mean Entropy
MNIST Diffusion + TreeVAE $\times$ $1.24$
TreeDiffusion $\checkmark$ 0.33
Fashion Diffusion + TreeVAE $\times$ $0.66$
TreeDiffusion $\checkmark$ 0.65
CIFAR10 Diffusion + TreeVAE $\times$ $1.12$
TreeDiffusion $\checkmark$ 0.93
CUBICC Diffusion + TreeVAE $\times$ 0.07
TreeDiffusion $\checkmark$ $0.20$
@@
For most datasets, the conditional model exhibits lower mean entropy, indicating that cluster conditioning indeed helps guide the model to generate more distinct and representative images for each leaf. However, for the CUBICC dataset, we observe that the mean entropy is lower for the cluster-unconditional model. This is because the classifier tends to predict all images into a single class, a result of model degeneration, where it primarily generates images for only a few classes. Figure˜6 visually presents the leaf generations for one sample of these models alongside the underlying TreeVAE generations. It can be observed that both the cluster-unconditional and conditional models exhibit a significant improvement in image quality. However, the images in the cluster-conditional model are more diverse, demonstrating greater adaptability for each cluster. Notably, across all models, the leaf-specific images share common properties, such as background color and overall shape, sampled at the root while varying in cluster-specific features from leaf to leaf within each model.
@@ 4.3 Ablation study on conditioning information
Finally, we conduct an ablation study to evaluate the impact of different conditioning signals on the generative performance. Specifically, we compare three types of conditioning: (i) hierarchical clustering signals derived from the latent embeddings of the selected cluster path $\bm{z}_{\mathcal{P}_{l}}$, (ii) flat clustering signals, including leaf assignment $l$, leaf embedding $\bm{z}_{l}$, or both, and (iii) an unconditioned setting where the diffusion model does not utilize any latent cluster representations from TreeVAE. Additionally, we examine the effect of using the TreeVAE leaf reconstruction $\hat{{\bm{x}}}_{0}^{(l)}$ as the starting point for the denoising process in the second-stage diffusion model. The results, outlined in Table 3, show the FID score calculated from $10,000$ samples generated using $100$ DDIM steps, averaged over $10$ seeds. Note that the first row in the table represents the TreeVAE + Diffusion model from the previous experiments, whereas the last row corresponds to the proposed TreeDiffusion method.
Table 3: Effect of conditioning signals on generative performance for CIFAR-10. FID scores for $10,000$ samples (lower is better) computed across $10$ random model initializations.
Conditioning Type $\hat{{\bm{x}}}_{0}^{(l)}$ $l$ $\bm{z}_{l}$ $\bm{z}_{\mathcal{P}_{l}}$ FID $\downarrow$
No Cluster Conditioning $19.7\,\text{{\footnotesize$\pm\,0.2$}}$
Flat Clustering-Based Conditioning $19.1\,\text{{\footnotesize$\pm\,0.3$}}$
$18.9\,\text{{\footnotesize$\pm\,0.3$}}$
$19.2\,\text{{\footnotesize$\pm\,0.2$}}$
$19.1\,\text{{\footnotesize$\pm\,0.5$}}$
Hierarchical Clustering-Based Conditioning $18.2\,\text{{\footnotesize$\pm\,0.3$}}$
@@
The findings suggest that incorporating latent leaf information — whether through leaf assignment, leaf embedding, or both — significantly improves generative performance compared to relying solely on leaf reconstructions. This highlights the added benefit of conditioning on flat clustering information. Furthermore, conditioning on the full path $\bm{z}_{\mathcal{P}_{l}}$, which integrates all embeddings and intermediate node assignments from the root to the leaf, leads to an even greater performance boost. This underscores the effectiveness of hierarchical clustering information beyond flat clustering. As a result, harnessing $\bm{z}_{\mathcal{P}_{l}}$ from the hierarchical structure not only produces more structured generations, as illustrated in Figure˜4, but also enhances the generative performance of generative clustering models. Notably, when conditioning on the full path, the model performs better without relying on TreeVAE reconstructions. Instead, the conditional diffusion model generates new images from scratch, guided solely by the latent information.
@@ 5 Conclusion
In this work, we present TreeDiffusion, a novel approach to integrate hierarchical clustering into diffusion models. By enhancing TreeVAE with a Denoising Diffusion Implicit Model conditioned on latent hierarchical representations, we propose a model capable of generating distinct, high-quality images that faithfully represent their respective data clusters. This approach not only improves the visual fidelity of the generated images but also facilitates cluster visualization. TreeDiffusion offers a robust framework that bridges the gap between clustering and generative performance, thereby expanding the potential applications of generative models in areas requiring detailed and more interpretable visual data interpretation.
{credits}
5.0.1 Acknowledgements
@@
Jorge da Silva Gonçalves is supported by the grant #2021-911 of the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). Laura Manduchi is supported by the SDSC PhD Fellowship #1-001568-037. Moritz Vandenhirtz is supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00047.
Appendix
Appendix 0.A Detailed Formulation of TreeVAE